OpenStudy (anonymous):

By factoring how do you solve 3x^2/3-11x^1/3-4=0? My teacher doesn't do substitution. please and thank you.

OpenStudy (cwrw238):

(3x^1/3 + 1)(x^1/3 - 4 ) = 0

OpenStudy (cwrw238):

can you continue?

OpenStudy (cwrw238):

3x^1/3 + 1 = 0 x^(1/3) = -1/3 x^1/3 - 4 = 0 x^1/3 = 4

OpenStudy (anonymous):

let me see...

OpenStudy (cwrw238):

hint: x = x(1/3)^3

OpenStudy (anonymous):

(-1/3)^3 or (1/3)^3?

OpenStudy (cwrw238):

sorry - i didn't write it corruptly (x^(1/3)^ 3 = x^(1/3 * 3) = x^1 = x or in other words you have the cube root of and you cube it - that gives you x so if you cube both sides of the equations you get the value of x e g (x^(1/3)^ 3 =( -1/3)^3 so x = - 1/3^3 = -1/27 thats one of the roots

OpenStudy (cwrw238):

* correctly

OpenStudy (anonymous):

Thank you so much.

OpenStudy (cwrw238):

note x^(1/3) is same as cube root x

OpenStudy (cwrw238):

so to find other root you do same thing with x^1/3 = 4

OpenStudy (cwrw238):

yw

OpenStudy (anonymous):

Thank you again. But do you divide x^1/3=4? Like 4/1=4.

OpenStudy (cwrw238):

just cube both sides of th equation as we did to the first one (x^(1/3)^3 = x 4^3 = 4*4*4 = 64 x = 64

OpenStudy (cwrw238):

i suggest you review the laws of indices x^a + x^b = x^(a+b) x^a / x^b = x^(a-b) (x^a)^b = x^(a*b) = x^ ab

OpenStudy (cwrw238):

also